OpenStudy (anonymous):

y varies inversely as twice x. When x = 3, y = 6. Find y when x = 2. y = 3 y = 6 y = 9 y = 12

OpenStudy (primeralph):

@marissalovescats Nooooo, don;t answer it yet.

OpenStudy (marissalovescats):

Okay... I was just going to give her the equation for inverse variation.

OpenStudy (primeralph):

Okay, let's do it @marissalovescats

OpenStudy (marissalovescats):

Well... It's y=k/x.... I think haha. I can't remember everything. So you plug in y and x to find K then use the k you get and the next x given to fond y.

OpenStudy (anonymous):

okay can you walk me through it cause i'm new at this stuff.

OpenStudy (marissalovescats):

Is that the inverse variation equation? Or is it y=kx? Lol

OpenStudy (primeralph):

y = k/x^n

OpenStudy (marissalovescats):

It's y=kx cuz I get an answer that's actually an answer. There are no exponents in inverse variation I know that for sure @primeralph

OpenStudy (primeralph):

Inverse variation? That's k/x^n

OpenStudy (marissalovescats):

I don't remember the equation but I aced Algebra 2 and I really dont think theres an exponent. Direct variation is y=kx Inverse= xy=k or y=k/x Joint= z=kxy..

OpenStudy (anonymous):

Since, y varies inversely as twice x therefore \[y \alpha \frac{1}{2x} \] i.e. \[y =k \frac{1}{2x} \rightarrow y =\frac{K}{2x}\] Where K is any constant. Now at x = 3, y = 6, K=? \[y =\frac{K}{2x} \rightarrow 6 =\frac{K}{2\times3} \rightarrow K=36\] Thus our eq of Variation becomes \[ y =\frac{36}{2x} \rightarrow y =\frac{18}{x}\] Now at x=2 y=? \[ y =\frac{18}{x} \rightarrow y =\frac{18}{2} \rightarrow y=9\]

OpenStudy (anonymous):

Thus y = 9 is the correct answer.

OpenStudy (primeralph):

@marissalovescats When you do more math, you'll begin to see the n.